Solve the following system of equations.
x+3y=4
x=6-3y

Answer by oscargut(2103)

x+3y=4
x=6-3y
Then:
6-3y+3y = 4
6 = 4 (not valid)
Answer: No solution , system is inconsistent
I can help you with all of your questions at: mthman@gmail.com
Thanks